OpenStudy (anonymous):

What is the simplified radical form of this exponent: 7^-(2/3)

OpenStudy (anonymous):

\(\frac{1}{\sqrt[3]{7^2}}\)

OpenStudy (anonymous):

your exponent is \(-\frac{2}{3}\) which has 1) a numerator 2 2) a denominator 3 3) a minus sign they mean 1) square 2) take the cubed root 3) take the reciprocal (flip it)

OpenStudy (anonymous):

can you should me how that would go? @satellite73

OpenStudy (anonymous):

you mean as a number? you need a calculator for this one

OpenStudy (anonymous):

Would it just be:\[1 \over \sqrt[3]{49}\]

OpenStudy (anonymous):

if you had \[8^{-\frac{2}{3}}\] we could say 1) cube root of 8 is 2 2) 2 squared is 4 3) the reciprocal of 4 is \(\frac{1}{4}\) so \[8^{-\frac{2}{3}}=\frac{1}{4}\]

OpenStudy (anonymous):

yes, you could write \[\frac{1}{\sqrt[3]{49}}\]

OpenStudy (anonymous):

but isn't it that you can't have a square root in the denominator?

OpenStudy (anonymous):

@satellite73